# Non-isothermal diffuse interface model for phase transition and interface evolution

### v1

@@ 2.5 Summary of non-isothermal systems
Now, we can summarize three non-isothermal diffuse interface models derived in this section by combining the results from Theorems 2.1, 2.4 and 2.7.
The equations are defined on the domain $(x,t)\in\Omega\times(0,T)$ and they shall be supplemented by a proper boundary conditions on $\partial\Omega\times(0,T)$, say the non-flux boundary condition, given by $\nabla\phi\cdot n=0,~{}~{}\nabla\theta\cdot n=0,~{}~{}u\cdot n=0,$.
@@
Now, assume that the assumption (A1) of Section 2.2 holds true. Then, the following system on $(\phi(x,t),\theta(x,t))$ satisfies both the first and second laws of thermodynamics
@@
$\begin{cases}&\partial_{t}\phi+u\cdot\nabla\phi=0,\quad-(u-v)\cdot\nabla\phi\nabla\phi\eta(\phi,\theta)=\mu\nabla\phi,\\ &\partial_{t}v+(v\cdot\nabla)v+\nabla p=\lambda\Delta v-\mu\nabla\phi,\quad\nabla\cdot v=0,\\ &\theta(\partial_{t}s+\nabla\cdot(sv))=\theta\nabla\cdot\left(\frac{\kappa(\phi,\theta)\nabla\theta}{\theta}\right)+\theta\Delta^{*},\\ &\theta\Delta^{*}=|(u-v)\nabla\phi|^{2}\eta(\phi,\theta)+\lambda|\nabla v|^{2}+\frac{\kappa(\phi,\theta)|\nabla\theta|^{2}}{\theta},\\ \end{cases}$
@@
$\partial_{t}v+(v\cdot\nabla)v+\nabla p=\lambda\Delta v-\mu\nabla\phi,\quad\nabla\cdot v=0,$
@@
(2.19) $\psi(\phi,\theta)=\theta(\frac{\epsilon}{2}|\nabla\phi|^{2}+\frac{1}{\epsilon}W(\varphi)+\frac{\chi}{2\sigma}(\theta-\theta_{c})\varphi)+c_{v}\theta(1-\log\theta),$
@@
(2.20) $s=-\psi_{\theta}=-\frac{\chi}{2\sigma}(2\theta-\theta_{c})\varphi+c_{v}\log\theta$
@@
(2.21) $e=\psi+\theta s=c_{v}\theta-\frac{\chi}{2\sigma}\theta_{c}\varphi.$
This choice of free energy leads to a heat equation that similar to that in Caginalp’s model [11, 12, 14, 13], in which the temperature equation is linearly depend on $\theta$ and $\varphi$.
3 The Stefan problem as a reduced model of the non-isothermal Allen-Cahn equation
In this section, we show that the sharp interface limit of the non-isothermal Allen-Cahn model (model with the kinematics assumption (A2) on the temperature) leads to a two-phase Stefan problem. We utilize the method of matched asymptotics, which has been successfully applied to various phase field models [6, 28, 27]. The idea behind a matched asymptotic expansion is to construct an approximation in the form of an asymptotic series that is obtained by splitting the domain into two parts; the inner problem and the outer problem, named for their relationship to the transition layer; and treating each part of the domain as a separate perturbation problem. The outer and inner solutions are then combined through a matching process such that an approximate solution for the whole domain is obtained.
We assume that $W(\phi)$ is the standard double-well potential and $g(\phi)=-\frac{1}{3}\phi^{3}+\phi$. In order to obtain the Stefan problem as the sharp interface limit, it is crucial to use the following scaling of the parameters: $\chi(\epsilon)=1$ and $\eta(\phi,\theta)=\epsilon$. The final non-isothermal Allen-Cahn model considered is
@@ 3.1 Outer expansion
Then, the leading order in (3.1) gives
@@
The solutions to (3.5) correspond to the minima of $W$ and are $\phi=0=\pm 1$. Thus, in the limit there are two regions corresponding to each minima defined by
@@
and the interface given by
@@
In the next order we obtain from (3.1) that on the interface $\Gamma_{0}$ we have that $\theta_{0}=\theta_{c}$.
For (3.2) we observe that $\partial_{t}\phi_{0}=0$ in the two pure domains $\Omega_{-}$ and $\Omega_{+}$. Moreover, by the smoothness of $k(\phi)$ we have that $k_{1}=k(\phi=1)$ in $\Omega_{-}$ and $k_{-1}=k(\phi=-1)$ in $\Omega_{+}$, respectively. Then, to the zeroth order we obtain
(3.8) $c_{v}\partial_{t}\theta=k_{\pm 1}\Delta\theta,\quad\textnormal{in }\Omega_{\pm}.$
@@ 3.2 Inner expansion and matching conditions
We now make an expansion in the interfacial region where the transition between the two phases takes place.
We denote $\Gamma_{0}$ the limiting hypersurface of the zero level sets of $\phi_{\varepsilon}$. And in order to study the limiting behaviour in these parts of $\Omega$ we introduce a new coordinate system. To this end, let $d(x)$ be the signed distance function to $\Gamma_{0}$ and introduce the new interface variable by $z=\frac{d}{\varepsilon}$, where we use the convention that $d(x)>0$ in $\Omega_{+}$ and $d(x)<0$ in $\Omega_{-}$. Additionally, let $g(t,s)$ denote a parametrization of $\Gamma_{0}$ by the arc length $s$, and let $\nu$ denote the unit normal of $\Gamma_{0}$ pointing into the domain $\Omega_{-}$. Then, in a tubular neighborhood of $\Gamma_{0}$, for a sufficiently smooth function $f(x)$, we have that
$f(x)=f(g(t,s)+\varepsilon z\nu(g(t,s)))=:F(t,s,z).$
In this new $(t,s,z)$-coordinate system the following rules of differentiation apply
@@
[…]
where $\mathcal{V}$ is the normal velocity of the interface $\Gamma_{0}$, $\nabla_{\Gamma_{0}}$ denotes the gradient on $\Gamma_{0}$ and $\kappa$ is the mean curvature of $\Gamma_{0}$. Here, h.o.t. denotes the higher order terms with respect to $\varepsilon$.
Next, we denote the functions $\phi_{\varepsilon},\theta_{\varepsilon}$ in the new coordinate system by $\Phi_{\varepsilon},\Theta_{\varepsilon}$, respectively. Then, we further assume that for $F_{\varepsilon}\in\{\Phi_{\varepsilon},\Theta_{\varepsilon}\}$ we have the following inner expansion
$F_{\varepsilon}(s,z)=F_{0}(s,z)+\varepsilon F_{1}(s,z)+\mathcal{O}(\varepsilon^{2}).$
As $\Gamma_{0}$ is the limit of the zero level sets of $\phi_{\varepsilon}$ this implies that
@@
Moreover, we assume that
@@
In order to match the two expansions we employ the following matching conditions
@@
Now, we have all the necessary tools at hand to study the inner expansion.
The leading order, i.e. the terms of order $\mathcal{O}(\epsilon^{-2})$, for (3.1) yield
@@
Using (3.9), we note that $\Phi_{0}$ can be chosen to be independent of $s$ and $t$, i.e. $\Phi_{0}$ is only a function of $z$. Thus $\Phi_{0}$ solves the ODE
@@
Additionally, we can multiply (3.15) by $\Phi_{0}^{\prime}(z)$ and obtain after integrating and applying the matching conditions the so-called equipartition of energy
(3.18) $\frac{1}{2}|\Phi_{0}^{\prime}(z)|^{2}=W(\Phi_{0}(z))=\frac{1}{4}\big{(}\Phi_{0}(z)^{2}-1\big{)}^{2}~{}~{}\textnormal{for all }|z|<\infty.$
Moreover, we can compute
$\int_{-\infty}^{\infty}W(\Phi_{0}(z))\,\textnormal{d}z=\frac{\sqrt{2}}{3}.$
Next, we consider the leading order in (3.2) and we obtain that
$\partial_{z}\big{(}k(\Phi_{0}(z))\partial_{z}\Theta_{0}\big{)}$ $=0.$
Integrating the equation once and using matching condition (3.12) yields that
(3.19) $k(\Phi_{0}(z))\partial_{z}\Theta_{0}=0.$
After integrating again and using the matching condition (3.11) together with the fact that $0<k(\Phi)\leq C$ we have
(3.20) $[\theta_{0}]_{+}^{-}=0,$
which corresponds to $\Theta_{0}(t,s,z)=\theta_{c}$, i.e. $\theta(x)=\theta_{c}$ for all $x\in\Gamma_{0}$.
@@
$=\big{[}W^{\prime}(\Phi_{0})\Phi_{1}-\partial_{z}\Phi_{1}\Phi_{0}^{\prime}\big{]}_{-\infty}^{\infty}+\int_{-\infty}^{\infty}\partial_{z}\big{(}W^{\prime}(\Phi_{0})-\Phi_{0}^{\prime\prime}\big{)}\,\textnormal{d}z$
$=0,$
@@
Lastly, for (3.2) we have
$-c_{v}\mathcal{V}\partial_{z}\Theta_{1}$ $=k(\Phi)\partial_{zz}\Theta_{1}-\kappa k(\phi)\partial_{z}\Theta_{0}-\frac{\lambda}{\theta_{c}}(\Phi_{0}^{2}-1)\Theta_{0}\mathcal{V}\Phi_{0}^{\prime}+|\mathcal{V}\Phi_{0}^{\prime}|^{2}.$
Integrating with respect to $z$ from $\infty$- ∞ to $\infty$ and applying the matching conditions (3.11)-(3.13) yields
(3.23) $\big{[}\nabla\theta_{0}\cdot\nu\big{]}_{-}^{+}=\mathcal{V}(\lambda\sqrt{2}-\mathcal{V})\frac{2\sqrt{2}}{3}.$
@@ 3.3 Sharp interface limit
Combining the inner and outer expansion of the asymptotic limit yields the following sharp interface system
(3.24) $\begin{split}\partial_{t}\theta&=\kappa_{+1}\Delta\theta~{}\textnormal{ in }\Omega_{+},\\ \partial_{t}\theta&=\kappa_{-1}\Delta\theta~{}\textnormal{ in }\Omega_{-},\\ \theta&=\theta_{c}~{}\textnormal{ on }\Gamma_{0},\\ \big{[}\nabla\theta_{0}\cdot\nu\big{]}_{-}^{+}&=\mathcal{V}(\lambda\sqrt{2}-\mathcal{V})\frac{2\sqrt{2}}{3}~{}\textnormal{ on }\Gamma_{0},\end{split}$
where the motion of the interface $\Gamma_{0}$ is given by
(3.25) $\begin{split}\mathcal{V}&=\kappa.\end{split}$
@@
We note that the limit system is indeed a two-phase Stefan-like problem. Hence, our result is in accordance with the observations by Caginalp.
Remark 3.3.
We compare the limit problem obtained from the non-isothermal Allen-Cahn system with the classical two-phase Stefan problem, which reads as follows
@@
(3.28) $\mathcal{V}(L_{l}-L_{s})$ $=-\big{[}\frac{k_{s/l}}{\rho_{s/l}}\partial_{\nu}\theta\big{]},\quad\textnormal{on }\Gamma_{0}$

### v2

@@ 2.5 Summary of non-isothermal systems
Now, we can summarize three non-isothermal diffuse interface models derived in this section by combining the results from Theorems 2.1, 2.4 and 2.7. The equations are defined on the domain $(x,t)\in\Omega\times(0,T)$ and they shall be supplemented by a proper boundary conditions on $\partial\Omega\times(0,T)$, say the non-flux boundary condition, given by $\nabla\phi\cdot n=0,~~\nabla\theta\cdot n=0,~~u\cdot n=0$.
@@
Assuming that the assumption (A1) of Section 2.2 holds true, the following system on $(\phi(x,t),\theta(x,t))$ satisfies both the first and second laws of thermodynamics:
@@
$\begin{cases}&\partial_{t}\phi+u\cdot\nabla\phi=0,\quad-(u-v)\cdot\nabla\phi\nabla\phi\eta(\phi,\theta)=\mu\nabla\phi,\\ &\partial_{t}v+(v\cdot\nabla)v+\nabla p=\lambda\Delta v-\mu\nabla\phi-s\nabla\theta,\quad\nabla\cdot v=0,\\ &\theta(\partial_{t}s+\nabla\cdot(sv))=\theta\nabla\cdot\left(\frac{\kappa(\phi,\theta)\nabla\theta}{\theta}\right)+\theta\Delta^{*},\\ &\theta\Delta^{*}=|(u-v)\nabla\phi|^{2}\eta(\phi,\theta)+\lambda|\nabla v|^{2}+\frac{\kappa(\phi,\theta)|\nabla\theta|^{2}}{\theta},\\ \end{cases}$
@@
$\partial_{t}v+(v\cdot\nabla)v+\nabla p=\lambda\Delta v-\mu\nabla\phi-(\chi(\epsilon)\frac{\lambda}{\theta_{c}}g(\phi)+c_{v}\log\theta)\nabla\theta,\quad\nabla\cdot v=0,$
@@
(2.20) $\psi(\phi,\theta)=\theta(\frac{\epsilon}{2}|\nabla\phi|^{2}+\frac{1}{\epsilon}W(\phi)+\frac{\chi}{2\sigma}(\theta-\theta_{c})\phi)+c_{v}\theta(1-\log\theta),$
@@
(2.21) $s=-\psi_{\theta}=-\frac{\chi}{2\sigma}(2\theta-\theta_{c})\phi+c_{v}\log\theta$
@@
(2.22) $e=\psi+\theta s=c_{v}\theta-\frac{\chi}{2\sigma}\theta_{c}\phi.$
This choice of free energy leads to a heat equation that similar to that in Caginalp’s model [11, 12, 14, 13], in which the temperature equation is linearly depend on $\theta$ and $\phi$.
3 The Stefan type problem as a reduced model of the non-isothermal Allen-Cahn equation
In this section, we show that the sharp interface limit of the non-isothermal Allen-Cahn model (model with the kinematics assumption (A2) on the temperature) leads to a two-phase Stefan type problem. We utilize the method of matched asymptotics, which has been successfully applied to various phase field models [6, 28, 27]. The idea behind a matched asymptotic expansion is to construct an approximation in the form of an asymptotic series that is obtained by splitting the domain into two parts; the inner problem and the outer problem, named for their relationship to the transition layer; and treating each part of the domain as a separate perturbation problem. The outer and inner solutions are then combined through a matching process such that an approximate solution for the whole domain is obtained.
We assume that $W(\phi)$ is the standard double-well potential and $g(\phi)=-\frac{1}{3}\phi^{3}+\phi$. In order to obtain a Stefan type problem as the sharp interface limit, it is crucial to use the following scaling of the parameters: $\chi(\epsilon)=1$ and $\eta(\phi,\theta)=\epsilon$. The final non-isothermal Allen-Cahn model considered is
@@ 3.1 Outer expansion
Substituting this expansion into the phase-field equation (3.1), the leading-order term at $\mathcal{O}(\varepsilon^{-2})$ yields
@@
The solutions to (3.5) are $\phi=0$ and $\phi=\pm 1$. Among these, $\phi=\pm 1$ correspond to the minima of the potential $W$, while $\phi=0$ is a saddle point. Thus, in the sharp-interface limit, the domain separates into two bulk regions corresponding to the pure phases:
@@
and the interface separating them is defined by
@@
At the next order, $\mathcal{O}(\varepsilon^{-1})$, equation (3.1) gives the interfacial condition:
(3.8) $\theta_{0}(x)=\theta_{c}~~\text{for }x\in\Gamma_{0}\ ,$
which indicates that the temperature must match the critical value $\theta_{c}$ on the interface.
Now, consider the temperature equation (3.2). In the pure phase regions $\Omega_{-}$ and $\Omega_{+}$, we have $\partial_{t}\phi_{0}=0$. Moreover, by the smoothness of $k(\phi)$, we define the constant thermal conductivities in each phase as
$k_{+1}=k(1),\quad k_{-1}=k(-1)$
Then, at leading order $\mathcal{O}(1)$ in (3.2), we obtain the classical heat equation in each region:
(3.9) $c_{v}\partial_{t}\theta_{0}=k_{\pm 1}\Delta\theta_{0},\quad\textnormal{in }\Omega_{\pm}.$
@@ 3.2 Inner expansion and matching conditions
We now perform an inner expansion in the interfacial region, where the transition between the two phases takes place.
@@
$f(x,t)=f(g(t,s)+\varepsilon z\nu(t,s),t)=:F(t,s,z).$
In this new $(t,s,z)$-coordinate system, the following rules of differentiation hold:
$\partial_{t}f(x,t)$ $=-\frac{1}{\varepsilon}\mathcal{V}\partial_{z}F+\mathcal{O}(1),$
$\nabla_{x}f(x,t)$ $=\frac{1}{\varepsilon}\partial_{z}F\cdot\nu+\nabla_{\Gamma_{0}}F+\mathcal{O}(\epsilon),$
$\Delta_{x}f(x,t)$ $=\frac{1}{\varepsilon^{2}}\partial_{zz}F-\frac{1}{\varepsilon}\kappa\partial_{z}F+\mathcal{O}(1),$
where $\mathcal{V}$ is the normal velocity of the interface $\Gamma_{0}(t)$, $\nabla_{\Gamma_{0}}$ denotes the gradient on $\Gamma_{0}$ and $\kappa$ is the mean curvature of $\Gamma_{0}$.
@@
$\Phi_{\varepsilon}(t,s,z):=\phi_{\varepsilon}(x,t),\quad\Theta_{\varepsilon}(t,s,z):=\theta_{\varepsilon}(x,t),$
where $x=g(t,s)+\varepsilon z\,\nu(t,s)$. Then, we further assume that for $F_{\varepsilon}\in\{\Phi_{\varepsilon},\Theta_{\varepsilon}\}$, the following inner expansion holds:
$F_{\varepsilon}(t,s,z)=F_{0}(t,s,z)+\varepsilon F_{1}(t,s,z)+\mathcal{O}(\varepsilon^{2}).$
Since $\Gamma_{0}(t)$ is the limiting zero level set of $\phi_{\varepsilon}$, it follows that the leading-order of the phase field function satisfies
@@
Moreover, we impose the far-field conditions:
@@
In addition, since the interface $\Gamma_{0}$ corresponds to the level set of the critical temperature $\theta_{c}$, we have
(3.12) $\Theta_{0}(t,s,z=0)=\theta_{c}.$
In order to match the inner and outer expansions, we impose the following matching conditions:
@@
We now derive the leading-order equation from the phase-field model. At leading order, the terms of order $\mathcal{O}(\varepsilon^{-2})$ in the equation (3.1) yield the following equation
@@
Using (3.10), we note that $\Phi_{0}$ can be chosen to be independent of $s$ and $t$, i.e. $\Phi_{0}$ is only a function of $z$. Thus, $\Phi_{0}$ solves the ODE
@@
To derive the energy identity, we multiply both sides of (3.17) by $\Phi_{0}^{\prime}(z)$:
$\Phi_{0}^{\prime\prime}(z)\Phi_{0}^{\prime}(z)=W^{\prime}(\Phi_{0}(z))\Phi_{0}^{\prime}(z).$
Recognizing both sides as total derivatives, we integrate with respect to $z$:
$\frac{1}{2}\frac{d}{dz}|\Phi_{0}^{\prime}(z)|^{2}=\frac{d}{dz}W(\Phi_{0}(z))\quad\Rightarrow\quad\frac{1}{2}|\Phi_{0}^{\prime}(z)|^{2}=W(\Phi_{0}(z))+C,$
where $C$ is an integration constant. To determine $C$, we use the matching conditions: as $z\to\pm\infty$, we have $\Phi_{0}(z)\to\pm 1$, hence $W(\Phi_{0}(z))\to 0$, and $\Phi_{0}^{\prime}(z)\to 0$. Substituting into the identity gives
$\frac{1}{2}|\Phi_{0}^{\prime}(\pm\infty)|^{2}=W(\Phi_{0}(\pm\infty))+C\quad\Rightarrow\quad 0=0+C,$
so $C=0$. Therefore, we arrive at the pointwise equipartition of energy:
(3.20) $\frac{1}{2}|\Phi_{0}^{\prime}(z)|^{2}=W(\Phi_{0}(z))\quad\text{for all }z\in\mathbb{R}.$
Moreover, we have
$\int_{-\infty}^{\infty}W(\Phi_{0}(z))\,\mathrm{d}z=\int_{-\infty}^{\infty}\frac{1}{2}|\Phi_{0}^{\prime}(z)|^{2}\,\mathrm{d}z=\frac{\sqrt{2}}{3}.$
Next, we consider the leading order $\mathcal{O}(\varepsilon^{-2})$ in (3.2) and we obtain that
$\partial_{z}\big{(}k(\Phi_{0}(z))\partial_{z}\Theta_{0}\big{)}$ =0=0.= 0 .
Integrating the equation once we obtain that $k(\Phi_{0}(z))\partial_{z}\Theta_{0}=\text{const.}$. Using the matching condition (3.14) implies then that
(3.21) $k(\Phi_{0}(z))\partial_{z}\Theta_{0}=0$
and from the interface condition $\Theta_{0}(st,z=0)=\theta_{c}$ it follows that
(3.22) $\Theta_{0}(t,s,z)=\theta_{c}.$
@@
$=\big{[}W^{\prime}(\Phi_{0})\Phi_{1}-\partial_{z}\Phi_{1}\Phi_{0}^{\prime}\big{]}_{-\infty}^{\infty}+\int_{-\infty}^{\infty}\partial_{z}\big{(}W^{\prime}(\Phi_{0})-\Phi_{0}^{\prime\prime}\big{)}\,\textnormal{d}z=0,$
@@
Lastly, for equation (3.2) we have in the order $\mathcal{O}(\epsilon^{-1})$ that
$-c_{v}\mathcal{V}\partial_{z}\Theta_{0}$ $=\partial_{z}\big{(}k(\Phi_{0})\partial_{z}\Theta_{1}\big{)}-\kappa k(\phi)\partial_{z}\Theta_{0}-\frac{\lambda}{\theta_{c}}(\Phi_{0}^{2}-1)\Theta_{0}\mathcal{V}\Phi_{0}^{\prime}+|\mathcal{V}\Phi_{0}^{\prime}|^{2}.$
We can reduce this by applying (3.22) and obtain
$\partial_{z}\big{(}k(\Phi_{0})\partial_{z}\Theta_{1}\big{)})=\lambda(\Phi_{0}^{2}-1)\mathcal{V}\Phi_{0}^{\prime}-|\mathcal{V}\Phi_{0}^{\prime}|^{2}$
Integrating with respect to $z$ from $\infty$- ∞ to $\infty$ and applying the matching condition (3.15) together with (3.20) yields
(3.25) $\big{[}k(\phi)\nabla\theta_{0}\cdot\nu\big{]}_{-}^{+}=\mathcal{V}(\lambda\sqrt{2}-\mathcal{V})\frac{2\sqrt{2}}{3}.$
@@ 3.3 Sharp interface limit
By performing a matched asymptotic expansion of the original system, we formally derive the following sharp-interface limit:
(3.26a) $c_{v}\partial_{t}\theta$ $=k_{+1}\Delta\theta~\textnormal{ in }\Omega_{+},$
(3.26b) $c_{v}\partial_{t}\theta$ $=k_{-1}\Delta\theta~\textnormal{ in }\Omega_{-},$
(3.26c) $\big{[}\nabla\theta\cdot\nu\big{]}_{-}^{+}$ $=\mathcal{V}(\lambda\sqrt{2}-\mathcal{V})\frac{2\sqrt{2}}{3}~\textnormal{ on }\Gamma_{0},$
together with the Dirichlet boundary condition for the temperature at the interface:
(3.27) $\theta=\theta_{c}~\textnormal{ on }\Gamma_{0}.$
Here, $\mathcal{V}$ is the the normal velocity at the interface $\Gamma_{0}$. The formal asymptotic analysis suggests that the interface evolves according to motion by mean curvature:
$\mathcal{V}=\kappa\ ,$
where $\kappa$ is the mean curvature of $\Gamma_{0}$. However, a rigorous mathematical analysis is still required to establish the existence of such a special solution to the Stefan-type problem (3.26) with the Dirichlet boundary condition (3.27).
@@
We note that the limit system is indeed a two-phase Stefan type problem. Hence, our result is in accordance with the observations by Caginalp. Thus, we can compare the limit problem obtained from the non-isothermal Allen-Cahn system with the classical two-phase Stefan problem, which reads as follows
@@
(3.30) $\mathcal{V}(L_{l}-L_{s})$ $=-\left[\frac{k_{s/l}}{\rho_{s/l}}\partial_{\nu}\theta\right]_{-}^{+},\quad\textnormal{on}\Gamma_{0}$
@@
Remark 3.3.
We want to point out that the quadratic nonlinearity in the jump condition (3.26c) is due to the entropy production in the system and therefore a direct consequence of the consistency of the system with the laws of thermodynamic.